### v1

# MET-Bench: Multimodal Entity Tracking for Evaluating the Limitations of Vision-Language and Reasoning Models

Abstract
Entity tracking is a fundamental challenge in natural language understanding, requiring models to maintain coherent representations of entities. Previous work has benchmarked entity tracking performance in purely text-based tasks. We introduce MET-Bench, a multimodal entity tracking benchmark designed to evaluate the ability of vision-language models to track entity states across modalities. Using two structured domains, Chess and the Shell Game, we assess how effectively current models integrate textual and image-based state updates. Our findings reveal a significant performance gap between text-based and image-based tracking and that this performance gap stems from deficits in visual reasoning rather than perception. We further show that explicit text-based reasoning strategies improve performance, yet substantial limitations remain, especially in long-horizon multimodal scenarios. Our results highlight the need for improved multimodal representations and reasoning techniques to bridge the gap between textual and visual entity tracking.
MET-Bench: Multimodal Entity Tracking for Evaluating
the Limitations of Vision-Language and Reasoning Models
Vanya Cohen  and Raymond Mooney Department of Computer Science The University of Texas at Austin {vanya,mooney}@utexas.edu
1 Introduction
Natural language understanding requires the ability to track and update information about entities as they evolve through text. From coreference resolution (Hobbs, 1978; Lappin and Leass, 1994) and discourse processing to narrative comprehension, computational linguistics has long grappled with the challenge of maintaining coherent entity representations across textual contexts (Bunescu and Paşca, 2006; Schank and Abelson, 1977).
While significant progress has been made in tasks like coreference resolution and entity linking Liu et al. (2023); Papalampidi et al. (2022), the broader challenge of tracking entity states—understanding how entities change through sequences of actions or events—remains an open challenge (Fagnou et al., 2024; Kim and Schuster, 2023; Toshniwal et al., 2022). This limitation becomes particularly apparent in tasks requiring integration of information across multiple modalities, an increasingly important frontier in computational linguistics as language processing systems are asked to reason about content that combines text with other forms of communication like images and video.
Our work examines this challenge through the lens of multimodal entity state tracking, where changes to entity states must be understood from both textual descriptions and visual observations. This setting provides a natural extension to classical NLP problems like discourse processing and situated language understanding, while also connecting to emerging research in multimodal dialogue and human-AI interaction. We focus specifically on scenarios where language models must reason about world-state changes described through a combination of text and images. Consider the task of understanding assembly instructions that combine written steps with supporting diagrams: while text might specify "Attach panel A to the base using the provided screws," accompanying images show the precise alignment and orientation. Accurate language understanding in such contexts requires maintaining a coherent mental model that integrates both linguistic and visual information about how entities’ states evolve.
(a) Chess State
(b) Chess Action
(c) Shell Game State
(d) Shell Game Action
Figure 1: Illustration of the two domains used in this work. (a) An example Chess board state. (b) The chess move (action) rendered as an image. (c) An initial shell game state with the blue ball under a shell. (d) Image action representing shells at positions one and three being swapped.
To systematically evaluate models’ capabilities in this multimodal language understanding setting, we introduce two complementary benchmarks: multimodal Chess and the Shell Game. Through these domains, we assess how effectively current language models can track entity states when updates are conveyed through both text and images. Our analysis reveals substantial disparities in how models process text-based and image-based entity-state updates, highlighting fundamental limitations in their multimodal language understanding.
While Chess and the Shell Game provide structured testbeds for evaluating entity tracking, real-world applications often involve more ambiguous and dynamic environments. However, by isolating state-tracking performance in controlled settings, we establish a clear baseline for assessing multimodal reasoning, one that provides a straightforward means of evaluation and can scale in difficulty with minimal changes.
We hypothesize that current language models struggle with multimodal entity tracking not due to low-level perceptual failures but because they lack representations (learned or otherwise engineered) for maintaining entity coherence across sequential visual observations. This suggests a fundamental limitation in how these models integrate and update state representations from different modalities.
We make the following contributions:
  • We introduce the multimodal entity tracking benchmark (MET-Bench) that extends traditional NLP entity tracking evaluation to the multimodal setting for two domains: multimodal Chess and Shell Game.
  • We demonstrate that current models, despite strong performance on pure text tasks, struggle to maintain accurate entity representations when processing mixed text and image inputs.
  • Through probing experiments, we show that these limitations stem from higher-level reasoning challenges rather than low-level perception issues.
  • We evaluate various approaches to improving multimodal entity tracking, finding that techniques emphasizing explicit reasoning outperform traditional NLP methods like fine-tuning when generalizing to novel domains.
2 Background
We formulate the problem of multimodal entity tracking as a sequential state estimation task, where an agent must infer the final state of a system given an initial state and a series of observed actions. Formally, at each time step $t$, the environment is in state $\bm{S_{t}}$, and transitions occur according to an action sequence $\bm{A}=(\bm{a_{1}},\bm{a_{2}},\dots,\bm{a_{T}})$. The agent receives observations $\bm{A_{t}}$ corresponding to each action, which may be textual ($\bm{A_{t}^{\text{text}}}$) or visual ($\bm{A_{t}^{\text{image}}}$). The objective is to infer the final state,
$\bm{S_{T}}=f(\bm{S_{0}},\bm{A_{1}},\bm{A_{2}},\dots,\bm{A_{T}}),$
where $f$ is the function modeling entity state updates.
MET-Bench represents the initial and final states of each domain as text but evaluate the models’ ability to track entity state changes through images. This approach isolates the multimodal entity tracking challenge by ensuring that models begin and end with well-defined textual representations while processing state transitions visually. By doing so, we assess their capacity to maintain coherent entity representations across modalities while minimizing confounding errors from perceptual failures, which remain a known limitation of current vision-language models Sharma et al. (2024).
2.1 Chess Domain
Chess is a well-studied domain for testing entity tracking of deep learning models Toshniwal et al. (2022). The state $S_{t}$ represents an 8×8 board configuration expressed in Forsyth–Edwards Notation (FEN) notation, actions correspond to legal chess moves from real games, and action observations consist of either symbolic (UCI notation) or visual (board images) descriptions of moves. We likewise adopt chess as an entity tracking testbed where the task is to maintain a correct representation of the board state across a sequence of moves. This distinction foregrounds how well a model can integrate and track piece locations as they change over time in potentially complicated board configurations.
Utilizing real Chess games from the Millionbase dataset [^1] https://rebel13.nl/rebel13/rebel%2013.html used in Toshniwal et al. (2022), we generate sequences of states and actions (moves) using standard chess notation: Universal Chess Interface (UCI) for actions and FEN for board states.
  • Text-Encoded Moves: Each action is provided as a short UCI textual description (e.g., ‘‘e2e4’’ for moving a piece from e2 to e4).
  • Image-Encoded Moves: Each action is accompanied by a rendered image that serves as a visual representation of the move (see Fig. 1).
In both cases, the final output is the FEN-encoded location of each piece on the board after a sequence of moves. The dataset includes multiple game trajectories of varying length, capturing a variety of piece types and board states.
2.2 Shell Game Domain
The second domain in MET-Bench is the Shell Game, a classic demonstration of hidden-state tracking. A ball is placed under one of three cups (or shells), which are then swapped pairwise in succession. The goal is tracking which cup currently hides the ball as shells are swapped. The state $S_{t}$ tracks the hidden position of a ball under three shells, actions correspond to swaps between pairs of shells. Other works have explored shell-game-like domains with varying levels of added complexity (Li et al., 2021; Long et al., 2016; Kim and Schuster, 2023).
This domain has a simpler entity-state and action space than Chess. However, while many frontier models have been trained on UCI/FEN encoded chess games, the Shell Game is, to the best of our knowledge, not present in the training data of these models. There may however be analogous tasks in the pre-training data.
We simulate repeated Shell Game swaps to create a set of Shell Game trajectories. Swap actions are either:
  • Text-Encoded Swap: Denoted as ‘‘x swap y’’, where $\{x,y\}$ are in $\{1,2,3\}$.
  • Image-Encoded Swap: An image depicting the shells being swapped, with the ball visually hidden (see Fig. 1).
The ground truth entity state after the game finishes is a single number indicating the final shell position of the ball.
2.3 Models
We use MET-Bench to evaluate the limitations of frontier models, including vision-language models (VLMs) which accept images and text as input, and newer reasoning models like OpenAI’s o1 that are trained using reinforcement learning and utilize test-time search algorithms to improve their reasoning abilities on domains like mathematics and coding. A full list of models and their capabilities is shown in Appendix A, Table 5.
3 Methods
Role Messages
User You are a helpful assistant that tracks chess moves in a game and produces the final FEN. The initial state is:
rnbqkbnr/pppppppp/8/8/8/8/ PPPPPPPP/RNBQKBNR w KQkq - 0 1
Here are the moves played:
e2e4
e7e5
Now what is the final FEN? Only output the FEN.
Assistant rnbqkbnr/pppp1ppp/8/4p3/4P3/ 8/PPPP1PPP/RNBQKBNR w KQkq - 2 2
Figure 2: An example zero-shot user–assistant exchange in the Chess domain, showing the initial board state as FEN, two UCI moves (e2e4, e7e5) and the final state. For image actions, the UCI moves are replaced with their visual representations and a description of how to interpret these images. The FEN is line-broken for readability.
We utilize the standard chat-based schema exposed by current frontier models that consists of interleaved user-provided and assistant (model) provided messages. Figures 2 shows the prompting strategy used for the Chess domain. Similarly, Appendix A, 6 shows the prompting strategy used for the Shell Game domain. Text actions are represented using simple notation on which the models have been trained, UCI for Chess and a simple domain-specific-language for Shell Game.
In the case of image-action input, the text actions are replaced with their image-rendered versions as Base64 encoded PNG images and a text description of how to interpret the image-actions is provided. Fig. 1 shows the image representations used for the Chess and Shell Game actions. These image representations were created through visual-prompt engineering to maximize the classification accuracy of actions depicted. Various common image representations were explored including arrows, bounding boxes, and symbolic markers. The image depiction of the game actions is explained to the language model every time images are provided using the prompts in Appendix A, Figures 7 & 8.
4 Experiments
Chess Text Image
Baseline
Game Start 74.4 74.4
Zero-Shot
Claude 3.5 Sonnet 96.8 66.2
MiniMax-VL-01 86.3 65.8
Gemini-2.0-Flash 93.3 74.7
GPT-4o mini 68.3 60.6
GPT-4o 89.6 73.3
Few-Shot (n=5)
Claude 3.5 Sonnet 99.6 77.9
MiniMax-VL-01 88.0 75.9
Gemini-2.0-Flash 96.6 78.7
GPT-4o mini 75.1 74.9
GPT-4o 94.9 77.7
Chain-of-Thought
Claude 3.5 Sonnet 97.9 68.4
MiniMax-VL-01 62.7 58.0
Gemini-2.0-Flash 93.3 74.5
GPT-4o mini 72.0 69.6
GPT-4o 91.8 74.9
Reasoning
o1-mini 65.1 -
o3-mini 99.6 -
o1 98.2 83.5
(a) Chess with ten moves in each sequence.
Shell Game Text Image
Baseline
Random 33.3 33.3
Zero-Shot
Claude 3.5 Sonnet 34.4 36.2
MiniMax-VL-01 34.4 35.2
Gemini-2.0-Flash 30.0 33.4
GPT-4o mini 30.6 32.8
GPT-4o 36.0 32.4
Few-Shot (n=5)
Claude 3.5 Sonnet 34.0 30.6
MiniMax-VL-01 36.4 32.0
Gemini-2.0-Flash 37.0 31.4
GPT-4o mini 34.4 31.2
GPT-4o 37.2 31.0
Chain-of-Thought
Claude 3.5 Sonnet 97.4 94.2
MiniMax-VL-01 92.6 32.8
Gemini-2.0-Flash 76.8 33.8
GPT-4o mini 84.4 35.0
GPT-4o 99.8 84.2
Reasoning
o1-mini 99.8 -
o3-mini 100.0 -
o1 100.0 92.8
(b) Shell Game with five swap moves in each sequence.
Table 1: Entity tracking accuracy in Chess and Shell Game for text and image actions. In the Few-Shot setting $N=5$ in-context examples are used. Methods and models which employ explicit reasoning perform best (chain-of-thought and reasoning models).
We perform experiments across a wide range of models and settings to evaluate different aspects of frontier-model entity-tracking performance. For all experiments, the models are sampled with a temperature of zero.
4.1 Tracking Entities with Text and Image Actions
We evaluate difference in accuracy when tracking images from text and image actions in the zero-shot, few-shot, and chain-of-thought settings. These results are presented in Section 5.1.
Zero-shot
Chess
In the Chess domain, we evaluate on a set of $100$ games selected at random from the test set, each with a sequence length of ten actions. The model must predict the FEN string for the final state. If the FEN string contains syntax errors such that it cannot be parsed, the accuracy for that instance is zero. We report the per-square accuracy of the predicted board, that is the ratio of correctly predicted pieces (or absence of a piece) to the total number of board tiles. The ‘Game Start’ baseline is the accuracy of predicting the initial board configuration. After only ten actions, most of the board configuration remains unchanged, so this is a strong baseline.
Shell Game
We evaluate Shell Game using a set of $500$ games generated at random, each with a sequence length of five actions. The final state is a single number $n\in\{1,2,3\}$ that gives the position of the ball, and we measure the accuracy of predicting it. The naive baseline picks a position uniformly at random.
Few-Shot and Chain-of-Thought
In these settings, the evaluation and procedure remain largely unchanged from the zero-shot Chess. For the few-shot experiments, $N=5$ examples are selected at random from the training set and prepended to the test example. The few-shot examples have the same number of actions as the test-set examples. To evaluate the effect of chain-of-thought reasoning, we prompt the model to ‘think step by step before producing a final answer.’
4.2 Sequence-Length Variation
The sequence length of the actions is varied to quantify the effect of compounding errors on the models. These sequences range from zero to $100$ actions in both the image and text action modality. This serves to quantify the relative drop-off in performance among models in the zero-shot setting. These results are presented in Section 5.2.
4.3 Image-Action Classification
We perform an experiment to demonstrate that VLMs have the ability to accurately interpret the actions depicted in the image-action representations. Figures 7 & 8 in Appendix A show the prompting strategy used to get a VLM to predict (zero-shot) a text action from an image action. We perform this evaluation on a test set of $1000$ image depictions of actions for each game. The results are presented in Section 5.3.
4.4 Cascaded Inference
Using the text actions predicted from the images in the image-action classification task, we devise an ablation to test the effect of cascading (performing multimodal inference in two steps) in the zero-shot setting. Given a sequence of image-based observations $\bm{A_{t}^{\text{image}}}$, a vision-language model (VLM) first predicts the corresponding text-based action sequence $\hat{\bm{A}}^{\text{text}}=g(\bm{A}^{\text{image}})$, where $g$ maps images to text using the procedure in Section 4.3. The text actions are then used for zero-shot inference to estimate the final state as $\bm{S_{T}}=f(\bm{S_{0}},\hat{\bm{A}}^{\text{text}})$. This removes the need for the model to perform entity tracking directly in the image modality, isolating the effect of perceptual failures. These results are presented in Section 5.4.
4.5 Fine-Tuning Experiments
We fine-tune frontier OpenAI models using their fine-tuning API on a small training set. We evaluate these in the Chess and Shell Game domains using both image and text modalities. For Chess, we select a random set of $100$ training examples of sequence length ten. For Shell Game we use $20$ training examples of sequence length $3$. Results are available in Section A.2. Hyperparameters were determined through grid search and are available in Appendix A.2, Table 6.
4.6 Mixed-Modality Experiments
We perform an ablation where the image-action and text-action modalities are mixed. At each action step, one action modality is randomly selected with a probability varying from $100\%$ text actions to $100\%$ image actions. This serves to quantify the effect of forcing the model to reason over multiple input modalities simultaneously. These results are presented in Section A.1.
5 Results
5.1 Tracking in Text Outperforms Images
Chess
Table 1(a) reports the accuracy for text and image actions in the chess domain. Performance in the text modality is significantly better than in the image modality. In the zero-shot setting, the best-performing text model, Claude 3.5 Sonnet, achieves an impressive $96.8\%$ accuracy, while its image-based counterpart drops sharply to $66.2\%$. A similar trend holds across models, indicating that current models can perform entity tracking in text but fail to integrate visual updates as effectively.
Both in-context learning with few-shot prompting and chain-of-thought reasoning lead to performance improvements. In the image modality, the accuracy is not significantly greater than the naive baseline. Only the o1 reasoning model achieves an accuracy greater than $80\%$ in the image modality.
Shell Game
(a) Chess zero-shot accuracy with text actions.
(b) Chess zero-shot accuracy with image actions.
Figure 3: In the text action setting, the reasoning models, o1 and o3-mini, maintain the highest accuracy at longer sequence lengths. o1-mini and the other models begin to output invalid and inaccurate board representations. All models struggle to maintain accurate board representations in the image action setting, with o1 performing the best.
Table 1(b) reports the accuracy for text and image actions for Shell Game. The results follow a pattern similar to that of Chess, with text-based tracking outperforming image-based tracking. However, unlike in Chess, the performance in the zero-shot setting is close to random. But the best model, o1, attains accuracies of $100.0\%$ and $92.8\%$ for the text and image modalities, respectively.
Few-shot prompting provides only marginal improvements, but chain-of-thought gives large performance increases. GPT-4o’s accuracy jumps from $36.0\%$ to $99.8\%$ in text, and from $32.4\%$ to $84.2\%$ in image tracking. These results suggest that when guided to decompose the task step-by-step, models can reason more effectively using image inputs, a finding that complements the results of performing cascaded inference in Section 5.4.
5.2 Reasoning Aids Long Sequence Accuracy
Chess
Figure 3(a) plots model accuracy against increasing sequence lengths of text actions. The models are evaluated in the zero-shot setting for sequence lengths of zero to $100$ text actions. Reasoning models like o1 and o3-mini are able to handle longer sequence lengths with a smaller decrease in accuracy. However, o1-mini performs worse than the other models as it produces more invalid FEN board representations at longer sequence lengths. In contrast, the non-reasoning models experience sharp decreases in accuracy after only a few actions. Figure 3(b) plots model accuracy against increasing sequence lengths of image actions. The models are evaluated in the zero-shot setting for sequence lengths of zero to $20$ image actions. While the reasoning models attain higher accuracies, the performance differences are smaller than in the text-action setting.
Shell Game
In the text modality in Figure 4(a), the reasoning models o1, o1-mini, and o3-mini attain the highest accuracies. o1 performs perfectly at a sequence length of 50 actions, where the non-reasoning models’ performance is significantly degraded. In the image modality results in Figure 4(b), o1 performs better than the other models, but sees a rapid decrease in accuracy with sequence lengths longer than five actions. By $20$ actions, the performance of all models has degraded to random.
The superior performance of reasoning models like o1 suggests that structured inference mechanisms, such as test-time search or chain-of-thought, are crucial for maintaining coherent entity states over long sequences. This aligns with prior findings that models trained on structured reasoning tasks (e.g., mathematics, coding) develop stronger implicit state-tracking capabilities (Kim et al., 2024), whereas standard vision-language models struggle with entity persistence beyond short contexts.
5.3 Models Understand Image Actions
(a) Shell Game zero-shot accuracy with text actions.
(b) Shell Game zero-shot accuracy with image actions.
Figure 4: In the text action setting, the reasoning models o1, o1-mini, and o3-mini achieve the highest performance over long action sequences, but performance degrades for all models as sequence length increases. In the image action setting, o1 performs the best, but achieves an accuracy no better than guessing the starting state by $20$ actions.
Model Start End Overall
Chess
GPT-4o-mini 62.1 55.2 41.0
GPT-4o 97.3 97.0 94.5
Shell Game
GPT-4o-mini 100.0 100.0 100.0
GPT-4o 100.0 100.0 100.0
Table 2: Percent image-action classification accuracy for various models. We report the accuracy of predicting the action start, end, and overall/UCI action for both Chess and Shell Game on $1{,}000$ image actions.
Cascaded
Method Chess Shell
Baseline
Game Start 74.4 33.3
Zero-Shot
GPT-4o mini 63.0 28.0
GPT-4o 89.8 34.0
Table 3: Cascaded entity tracking accuracy for Chess and Shell (Image → Text). In cascaded inference, the model is first used to map each image action to the text representation of the action. Then model is prompted to perform the entity tracking task as in the text-action setting.
Table 2 shows the performance of GPT-4o and GPT-4o-mini on classifying the text action represented by each image action. For Chess, the action ‘start’ is the square of the piece being moved, and ‘end’ is the destination square. For Shell Game, the ‘start’ is the first shell to be swapped and ‘end’ is the second. ‘Overall’ is the accuracy of classifying the entire action (start and end) correctly. While GPT-4o-mini struggles to recognize actions in chess, GPT-4o achieves an accuracy of $94.5\%$. Both models attain perfect accuracy on the simpler Shell Game domain. This indicates that perception of the image-actions is not the fundamental limiting factor for effective entity tracking with image inputs.
A potential concern is whether the observed failures stem from poor image representation rather than reasoning deficiencies. However, our cascaded inference experiments (Section 5.4) demonstrate that when models first translate image actions into text, they achieve near-text-level accuracy. This suggests that models can correctly parse image-based actions but struggle with integrating them into coherent state updates, a limitation in reasoning rather than perception.
5.4 Cascading Matches Text-Only Tracking
Table 3 shows the accuracy of cascaded inference in the zero-shot setting for GPT-4o and GPT-4o-mini. In this setting, the image actions are first translated into text actions, and then run through the text-based entity tracking pipeline. The performance in the cascaded setting is similar to the text-action performance, showing that the model has the task-knowledge needed to perform entity tracking in both domains, but cannot reason effectively in the image modality.
6 Discussion
Our evaluation of frontier model performance on MET-Bench provides several insights into the current state and remaining challenges of multimodal entity tracking. We demonstrate a significant performance gap between text-based and image-based entity tracking across all evaluated models, with even state-of-the-art vision-language-reasoning models struggling to maintain accurate entity states when processing visual inputs. This disparity persists across both the Chess and Shell Game domains, suggesting a fundamental limitation in current architectures’ ability to reason about entity states through visual observations.
This finding is particularly noteworthy given that our image-action classification results (Table 2) demonstrate that models can accurately perceive and classify individual visual actions. The gap between perception and reasoning suggests that the challenge lies not in processing visual inputs, but in maintaining and updating coherent entity information across sequential visual observations.
Our cascaded inference experiments provide further evidence for this interpretation. When models first translate visual inputs to text before performing entity tracking, they achieve performance comparable to pure text-based tracking. This indicates that the models possess the relevant task knowledge and reasoning capabilities, but struggle to apply them directly in the visual domain.
Further, the effectiveness of chain-of-thought prompting, particularly in the Shell Game domain where it improved GPT-4o’s accuracy from $36.0\%$ to $99.8\%$ for text and $32.4\%$ to $84.2\%$ for images, highlights the importance of explicit reasoning for entity tracking. This improvement indicates that current models can perform complex entity tracking when guided to decompose the task into smaller steps, even in novel domains not present in their training data. However, the fact that such prompting was necessary suggests that models do not implement robust tracking, particularly in multimodal settings. Lastly, the performance of specialized reasoning models like o1 and o3-mini on longer sequences demonstrates the potential of architectures explicitly trained for sequential reasoning to maintain coherent entity states despite the challenges of accumulating errors over extended sequences.
7 Related Work
Entity tracking has been extensively studied in textual domains, with a focus on probing and improving language models’ abilities to maintain representations of entity states. For instance, Toshniwal et al. (2022) evaluates chess as an entity tracking domain, employing fine-tuned models (Radford et al., 2019) to assess performance. Similarly, Kim and Schuster (2023) examine the impact of model size and fine-tuning on entity tracking in textual settings similar to our Shell Game domain. Tandon et al. (2020) construct a benchmark for understanding entity state changes in procedural texts. Shirai et al. (2022) construct the Visual Recipe Flow corpus and evaluate the ability of multimodal embedding models to properly sequence images depicting recipe states. In contrast, our work requires predicting entity state changes from actions specified in images and involves larger state spaces.
Several studies explore the implicit representations of entity states in language models. Li et al. (2021) and Long et al. (2016) use semantic probing to reveal that Transformer-based models (Vaswani et al., 2017) capture entity state representations implicitly during textual reasoning. Building on this, Prakash et al. (2024) demonstrate that fine-tuning language models for entity tracking tasks enhances pre-existing internal mechanisms rather than learning entirely new representations. Li et al. (2023) find that Transformers trained on Othello games form internal representations of the game state.
Efforts to improve textual entity tracking beyond domain-specific fine-tuning include Fagnou et al. (2024), which establishes theoretical limitations of the Transformer architecture in tracking entities. They propose a novel attention mechanism to enhance entity tracking in Transformers. Gupta and Durrett (2019) fine-tunes small Transformer-based models for tracking entity state in instructional texts. Kim et al. (2024) investigates how code pretraining improves language models’ abilities to track entities in text, while Yoneda et al. (2024) introduce Stalter, a prompting method designed to maintain accurate state representations in text-based robotics planning.
These works focus on entity tracking as a unimodal, text-based reasoning task. While unimodal approaches have achieved substantial progress, there remains a gap in evaluating models’ ability to integrate multimodal inputs for entity tracking. Our work extends these evaluations to the multimodal setting and quantifies the performance improvement of reasoning models for entity tracking.
8 Conclusion
Our findings suggest that the primary bottleneck in multimodal entity tracking is not visual recognition but sequential reasoning over visual updates. Unlike text-based representations, which align with the models’ training paradigms, visual updates require implicit state reconstruction—a task that current architectures do not perform reliably. Future work should explore the effect of additional visual-reasoning post-training, explicit memory structures, or hybrid symbolic representations to mitigate this gap. Additional research directions include investigating the role of entity tracking in world-modeling, narrative understanding, and expanding MET-Bench to include more complex domains beyond games. We believe addressing these challenges will be crucial for developing AI systems capable of robust reasoning for real-world tasks.
9 Limitations
Our benchmark is restricted to two synthetic domains—Chess and the Shell Game. While these domains are well-structured and offer a clear framework for assessing entity state tracking, they may not fully capture the complexity of real-world multimodal reasoning tasks that require entity tracking.
Our study highlights a substantial performance gap between text-based and image-based entity tracking, yet the exact causes of this disparity remain unclear. While our cascaded inference experiments suggest that models struggle to reason directly over visual updates rather than simply perceiving them, further investigation is needed to pinpoint the underlying source of the problem and whether it’s possible to explicitly ameliorate it.
Additionally, our fine-tuning experiments demonstrate that performance can be improved with domain-specific adaptation, but we do not investigate the trade-offs between fine-tuning and generalization across unseen multimodal datasets. Understanding whether these improvements transfer to novel multimodal reasoning tasks is an important avenue for future work.
Ethical Considerations
Vision-language models may inherit biases from their training data or other aspects of their development, leading to disparities in performance across different languages, cultures, or other categories that intersect with textual and visual representations. The evaluations in this work can be used to improve the capabilities of frontier models, which if deployed in contexts requiring multimodal entity tracking, could further expose end-users to the biases present in this work including our choice of game domains and language.
Acknowledgments
This paper and associated code were created with the assistance of ChatGPT for proofreading and debugging.
Appendix A Appendix
A.1 Models Struggle to Integrate Mixed Modalities
Figure 5: Chain-of-thought entity tracking accuracy for Chess and Shell Game with GPT-4o. The data splits range from 100% text-encoded actions to 100% image-encoded actions. The plot illustrates the change in accuracy as actions shift between modalities.
To examine how well models can integrate mixed-modality information, we evaluate performance as we vary the proportion of text and image-based action representations. As shown in Figure 5, for Chess performance degrades smoothly as the fraction of image actions increases, rather than exhibiting an abrupt collapse. However for Shell Game, the opposite is true and mixes of text and image actions are challenging for the model to reason over.
A.2 Fine-Tuning Improves Multimodal Entity Tracking
Fine-tuned Text Image
Chess
GPT-4o mini 89.2 -
GPT-4o 97.0 86.4
Shell Game (S=3)
GPT-4o mini 32.0 -
GPT-4o 74.0 32.0
Table 4: Fine-tuned model entity tracking accuracy for the Chess and Shell Game domains (Text actions and Image actions).GPT-4o-mini does not support image finetuning. A training set of $100$ action sequences of length ten were used for Chess and $20$ action sequences of length three and five for Shell Game.
Fine-tuning using the OpenAI fine-tuning API substantially improves model performance across both text and image modalities, as shown in Table 4. In Chess, fine-tuned models outperform even the strongest zero-shot reasoning models, achieving 97.0% accuracy in the text domain and a significant boost to 86.4% in the image domain. This suggests that even with a relatively small dataset, fine-tuning allows the model to learn entity tracking representations that generalize better in both modalities. Notably, fine-tuning leads to a larger improvement in the image modality than in the text modality. This reinforces the idea that pretrained models already encode strong textual reasoning capabilities, whereas multimodal reasoning requires additional adaptation.
In contrast, improvements on a simplified version of Shell Game with only three moves are minimal in case of image-encoded actions. The Shell Game task is not present in the training data and it’s harder for the model to generalize, even when exposed to a large fraction of the possible games of the given length. This may indicate that the Shell Game domain is simply too challenging for the model to learn in both the image and text settings from a limited number of examples. A more complex training curriculum involving fine-tuning over multiple game lengths may be required.
Model Name Image Reasoning
Claude 3.5 Sonnet
Anthropic (2024)
Gemini-2.0-Flash
Hassabis and Kavukcuoglu (2024)
GPT-4o mini
OpenAI (2024a)
GPT-4o
OpenAI (2024b)
Minimax-VL-01
MiniMax et al. (2025)
o1-mini
OpenAI (2024c)
o1
OpenAI (2024d)
o3-mini
OpenAI (2025)
Table 5: Comparison of capabilities of language models evaluated using the MET benchmark. All evaluated models support text input and output. The total API cost of experiments run is \$$2340.00$.
Configuration Epochs LRM Batch
Chess
   Text:
    GPT-4o-mini 3 1.8 1
    GPT-4o 3 2 1
   Image:
    GPT-4o 3 2 1
Shell Game
   Text:
    GPT-4o-mini 3 1.8 1
    GPT-4o 3 2 1
   Image:
    GPT-4o 5 2 1
Table 6: Hyperparameters used for fine-tuning across domains and modalities. The training epochs, learning rate multiplier (LRM), and batch size are reported.
Role Messages
User The shell game is a classic game where a ball is hidden under one of three shells. You are a helpful assistant that tracks the position of the ball. The ball starts under shell 2. Here are the moves played:
1 swap 3
2 swap 3
Now what is the final position of the ball? Only output the number 1, 2, or 3.
Assistant 3
Figure 6: An example zero-shot user–assistant exchange in the Shell Game domain, illustrating how the system tracks swaps to determine the ball’s final shell.
Role Messages
User You are a helpful assistant that interprets image-based actions in chess.
Here is an image representing a move:
[Image Input]
In UCI notation, what move does the arrow on the chessboard represent? The move is from the green square to the red square. (e.g., ‘e2e4’). Only output the move and nothing else.
Assistant e2e4
Figure 7: An example user–assistant exchange in the Chess domain, where the assistant identifies the move represented in the image.
Role Messages
User You are a helpful assistant that interprets image-based actions in the shell game.
Here is an image representing a swap:
[Image Input]
In shell game notation, which shells are being swapped in the image? Shells are labeled ‘1’, ‘2’, ‘3’ and the shells being swapped have their numbers highlighted in green. Only output a dash-separated pair like ‘1 swap 3’ and nothing else.
Assistant 1 swap 3
Figure 8: An example user–assistant exchange in the Shell Game domain, where the assistant identifies the shell swap represented in the image.
A.3 Models
The models evaluated using MET-Bench are listed in Table 5.
Minimax-VL-01
This model is released under the license: https://github.com/MiniMax-AI/MiniMax-01/blob/main/LICENSE. The model is 465 billion parameters and is trained on a “diverse [dataset] incorporating diverse sources including academic literature, books, web content, and programming code” and post-training dataset encompassing many multimodal and NLP tasks of 512 billion tokens (MiniMax et al., 2025).
A.3.1 Proprietary Models
These models have limited information about their training and development. Like Minimax-VL-01, these models are likely trained on diverse, web-scale corpora spanning many domains and tasks. We provide links to the current terms of their use.
Claude 3.5 Sonnet
Gemini-2.0-Flash
GPT-4o mini, GPT-4o, o1, o1-mini, o3-mini
A.4 Datasets
The Chess dataset is adapted from Toshniwal et al. (2022) which is adapted from the MillionBase dataset, available for download at https://rebel13.nl/rebel13/rebel%2013.html. To the best of our knowledge, no license or terms of use are currently listed for either the original MillionBase dataset or dataset of Toshniwal et al. (2022). Our usage of this dataset is consistent with the description of its use by Toshniwal et al. (2022).
MET-Bench is intended for evaluating and improving the ability of VLMs to perform entity tracking. It is released under the MIT License.

### v2

@@
 Abstract
-Entity tracking is a fundamental challenge in natural language understanding, requiring models to maintain coherent representations of entities.
+Entity state tracking is a necessary component of world modeling that requires maintaining coherent representations of entities over time.
 Previous work has benchmarked entity tracking performance in purely text-based tasks.
-We introduce MET-Bench, a multimodal entity tracking benchmark designed to evaluate the ability of vision-language models to track entity states across modalities.
-Using two structured domains, Chess and the Shell Game, we assess how effectively current models integrate textual and image-based state updates.
-Our findings reveal a significant performance gap between text-based and image-based tracking and that this performance gap stems from deficits in visual reasoning rather than perception.
-We further show that explicit text-based reasoning strategies improve performance, yet substantial limitations remain, especially in long-horizon multimodal scenarios.
+We introduce MET-Bench, a multimodal entity tracking benchmark designed to evaluate vision-language models’ ability to track entity states across modalities.
+Using two structured domains, we assess how effectively current models integrate textual and image-based state updates.
+Our findings reveal a significant performance gap between text-based and image-based entity tracking.
+We empirically show this discrepancy primarily stems from deficits in visual reasoning rather than perception.
+We further show that explicit text-based reasoning strategies improve performance, yet limitations remain in long-horizon multimodal tasks.
+We develop a reinforcement learning method to improve performance on MET-Bench.
+Applying our method to open-source VLMs achieves competitive performance with advanced closed models.
 Our results highlight the need for improved multimodal representations and reasoning techniques to bridge the gap between textual and visual entity tracking.
-MET-Bench: Multimodal Entity Tracking for Evaluating
-the Limitations of Vision-Language and Reasoning Models
-Vanya Cohen  and Raymond Mooney Department of Computer Science The University of Texas at Austin {vanya,mooney}@utexas.edu
+Machine Learning, ICML
 1 Introduction
-Natural language understanding requires the ability to track and update information about entities as they evolve through text.
-From coreference resolution (Hobbs, 1978; Lappin and Leass, 1994) and discourse processing to narrative comprehension, computational linguistics has long grappled with the challenge of maintaining coherent entity representations across textual contexts (Bunescu and Paşca, 2006; Schank and Abelson, 1977).
-While significant progress has been made in tasks like coreference resolution and entity linking Liu et al.
-(2023); Papalampidi et al.
-(2022), the broader challenge of tracking entity states—understanding how entities change through sequences of actions or events—remains an open challenge (Fagnou et al., 2024; Kim and Schuster, 2023; Toshniwal et al., 2022).
-This limitation becomes particularly apparent in tasks requiring integration of information across multiple modalities, an increasingly important frontier in computational linguistics as language processing systems are asked to reason about content that combines text with other forms of communication like images and video.
+Developing models that track and predict the latent state of the world has emerged as an important goal for artificial intelligence (worldmodels2024).
+One faculty of effective world models is entity state tracking: estimating the evolution of entities, their attributes and relations over time.
+Many problems require accurate entity tracking including robotic manipulation (yoneda2024statler; nvidia2025gr00tn1openfoundation), video–question‐answering (he2024mmworld), and computer-control agents (bishop2024latentstateestimationhelps; sager2025comprehensive).
+Success on these tasks requires maintaining a coherent estimate of how entity state changes from sequential, multimodal observations.
+Early progress on entity tracking emerged within text-only tasks.
+From coreference resolution (HOBBS1978311; Lappin and Leass, 1994) and discourse processing to narrative comprehension, computational linguistics has long grappled with the challenge of maintaining coherent entity representations across textual contexts (Bunescu and Paşca, 2006; Schank1977).
+While significant progress has been made in text-only tasks (liu2023brief; papalampidi2022towards), the broader challenge of tracking entity states—understanding how entities change through sequences of actions or events—remains an open challenge (Fagnou et al., 2024; Kim and Schuster, 2023; toshniwal2022chess).
+This limitation becomes particularly apparent in tasks requiring integration of information across multiple modalities, an increasingly important frontier in AI as systems are asked to reason about content that combines text with other modalities like images and video.
 Our work examines this challenge through the lens of multimodal entity state tracking, where changes to entity states must be understood from both textual descriptions and visual observations.
-This setting provides a natural extension to classical NLP problems like discourse processing and situated language understanding, while also connecting to emerging research in multimodal dialogue and human-AI interaction.
-We focus specifically on scenarios where language models must reason about world-state changes described through a combination of text and images.
-Consider the task of understanding assembly instructions that combine written steps with supporting diagrams: while text might specify "Attach panel A to the base using the provided screws," accompanying images show the precise alignment and orientation.
-Accurate language understanding in such contexts requires maintaining a coherent mental model that integrates both linguistic and visual information about how entities’ states evolve.
-(a) Chess State
-(b) Chess Action
-(c) Shell Game State
-(d) Shell Game Action
-Figure 1: Illustration of the two domains used in this work.
-(a) An example Chess board state.
-(b) The chess move (action) rendered as an image.
-(c) An initial shell game state with the blue ball under a shell.
-(d) Image action representing shells at positions one and three being swapped.
-To systematically evaluate models’ capabilities in this multimodal language understanding setting, we introduce two complementary benchmarks: multimodal Chess and the Shell Game.
+This setting provides a natural extension to classical NLP problems while also connecting to emerging research in multimodal dialogue, robotics, and human-AI interaction.
+We focus specifically on scenarios where multimodal models must reason about world-state changes described through a combination of text and images.
+Figure 1: In the multimodal entity tracking benchmark (MET-Bench), a vision-language model (VLM) predicts the final entity state from the initial state and actions which update the entity state.
+The initial entity state is provided as text and the actions are given as images or as text.
+The predicted final state is a text representation of the entity state.
+To systematically evaluate models’ capabilities in this multimodal setting, we introduce two complementary benchmarks: multimodal Chess and Shell Game.
 Through these domains, we assess how effectively current language models can track entity states when updates are conveyed through both text and images.
-Our analysis reveals substantial disparities in how models process text-based and image-based entity-state updates, highlighting fundamental limitations in their multimodal language understanding.
+Our analysis reveals substantial disparities in how models process text-based and image-based entity-state updates, highlighting fundamental limitations in their multimodal understanding.
 While Chess and the Shell Game provide structured testbeds for evaluating entity tracking, real-world applications often involve more ambiguous and dynamic environments.
-However, by isolating state-tracking performance in controlled settings, we establish a clear baseline for assessing multimodal reasoning, one that provides a straightforward means of evaluation and can scale in difficulty with minimal changes.
-We hypothesize that current language models struggle with multimodal entity tracking not due to low-level perceptual failures but because they lack representations (learned or otherwise engineered) for maintaining entity coherence across sequential visual observations.
-This suggests a fundamental limitation in how these models integrate and update state representations from different modalities.
-We make the following contributions:
+However, by isolating state-tracking performance in controlled settings, we establish a clear baseline for assessing multimodal reasoning, one that provides a straightforward means of evaluation where difficulty is easily controlled.
+Overall, this paper makes the following contributions:
   • We introduce the multimodal entity tracking benchmark (MET-Bench) that extends traditional NLP entity tracking evaluation to the multimodal setting for two domains: multimodal Chess and Shell Game.
   • We demonstrate that current models, despite strong performance on pure text tasks, struggle to maintain accurate entity representations when processing mixed text and image inputs.
-  • Through probing experiments, we show that these limitations stem from higher-level reasoning challenges rather than low-level perception issues.
-  • We evaluate various approaches to improving multimodal entity tracking, finding that techniques emphasizing explicit reasoning outperform traditional NLP methods like fine-tuning when generalizing to novel domains.
+  • Through experiments, we show that these limitations stem from higher-level reasoning challenges rather than simple perceptual issues.
+  • We apply reinforcement learning to improve the reasoning of open-source VLMs for entity tracking, and achieve performance competitive with advanced closed models.
+However, we find only limited transfer learning between modalities.
 2 Background
 We formulate the problem of multimodal entity tracking as a sequential state estimation task, where an agent must infer the final state of a system given an initial state and a series of observed actions.
-Formally, at each time step $t$, the environment is in state $\bm{S_{t}}$, and transitions occur according to an action sequence $\bm{A}=(\bm{a_{1}},\bm{a_{2}},\dots,\bm{a_{T}})$.
-The agent receives observations $\bm{A_{t}}$ corresponding to each action, which may be textual ($\bm{A_{t}^{\text{text}}}$) or visual ($\bm{A_{t}^{\text{image}}}$).
+Formally, at each time step $t$, the environment is in state $\mathbf{S}_{t}$, and transitions occur according to an action sequence $\mathbf{A}=(\mathbf{a}_{1},\mathbf{a}_{2},\dots,\mathbf{a}_{T})$.
+The agent receives $\mathbf{a}_{t}$ corresponding to each action, which may be textual $\mathbf{a}_{t}^{\text{text}}$ or visual $\mathbf{a}_{t}^{\text{image}}$.
 The objective is to infer the final state,
-$\bm{S_{T}}=f(\bm{S_{0}},\bm{A_{1}},\bm{A_{2}},\dots,\bm{A_{T}}),$
-where $f$ is the function modeling entity state updates.
-MET-Bench represents the initial and final states of each domain as text but evaluate the models’ ability to track entity state changes through images.
+$\mathbf{S}_{T}=\operatorname{f}(\mathbf{S}_{0},\mathbf{a}_{1},\mathbf{a}_{2},\dots,\mathbf{a}_{T}),$
+where $\operatorname{f}$ is the function modeling entity state updates.
+MET-Bench represents the initial and final states of each domain as text but also evaluates models’ ability to track entity state changes through images.
 This approach isolates the multimodal entity tracking challenge by ensuring that models begin and end with well-defined textual representations while processing state transitions visually.
-By doing so, we assess their capacity to maintain coherent entity representations across modalities while minimizing confounding errors from perceptual failures, which remain a known limitation of current vision-language models Sharma et al.
-(2024).
+This assesses the models’ capacity to maintain coherent entity representations across modalities while minimizing confounding errors from perceptual failures, which remain a limitation of current vision-language models (sharma2024vision).
 2.1 Chess Domain
-Chess is a well-studied domain for testing entity tracking of deep learning models Toshniwal et al.
-(2022).
+Chess is a well-studied domain for testing entity tracking of deep learning models (toshniwal2022chess).
 The state $S_{t}$ represents an 8×8 board configuration expressed in Forsyth–Edwards Notation (FEN) notation, actions correspond to legal chess moves from real games, and action observations consist of either symbolic (UCI notation) or visual (board images) descriptions of moves.
 We likewise adopt chess as an entity tracking testbed where the task is to maintain a correct representation of the board state across a sequence of moves.
-This distinction foregrounds how well a model can integrate and track piece locations as they change over time in potentially complicated board configurations.
-Utilizing real Chess games from the Millionbase dataset [^1] https://rebel13.nl/rebel13/rebel%2013.html used in Toshniwal et al.
-(2022), we generate sequences of states and actions (moves) using standard chess notation: Universal Chess Interface (UCI) for actions and FEN for board states.
-  • Text-Encoded Moves: Each action is provided as a short UCI textual description (e.g., ‘‘e2e4’’ for moving a piece from e2 to e4).
-  • Image-Encoded Moves: Each action is accompanied by a rendered image that serves as a visual representation of the move (see Fig. 1).
+Utilizing real games from Millionbase [^1] https://rebel13.nl/rebel13/rebel%2013.html (toshniwal2022chess), we generate sequences of states and actions (moves) using standard chess notation: Universal Chess Interface (UCI) for actions and FEN for board states.
+Text-Encoded Moves (each action is provided as a short UCI textual description, e.g., ‘‘e2e4’’ for moving a piece from e2 to e4); Image-Encoded Moves (each action is accompanied by a rendered image serving as a visual representation of the move; see Fig.
+1).
+The visual action depictions were created to maximize perceptual accuracy as in Table 4.
 In both cases, the final output is the FEN-encoded location of each piece on the board after a sequence of moves.
@@
 2.2 Shell Game Domain
-The second domain in MET-Bench is the Shell Game, a classic demonstration of hidden-state tracking.
+The second domain is Shell Game, a classic demonstration of hidden-state tracking.
 A ball is placed under one of three cups (or shells), which are then swapped pairwise in succession.
-The goal is tracking which cup currently hides the ball as shells are swapped.
+The goal is tracking which cup currently hides the ball.
 The state $S_{t}$ tracks the hidden position of a ball under three shells, actions correspond to swaps between pairs of shells.
 Other works have explored shell-game-like domains with varying levels of added complexity (Li et al., 2021; Long et al., 2016; Kim and Schuster, 2023).
-This domain has a simpler entity-state and action space than Chess.
-However, while many frontier models have been trained on UCI/FEN encoded chess games, the Shell Game is, to the best of our knowledge, not present in the training data of these models.
-There may however be analogous tasks in the pre-training data.
-We simulate repeated Shell Game swaps to create a set of Shell Game trajectories. Swap actions are either:
-  • Text-Encoded Swap: Denoted as ‘‘x swap y’’, where $\{x,y\}$ are in $\{1,2,3\}$.
-  • Image-Encoded Swap: An image depicting the shells being swapped, with the ball visually hidden (see Fig. 1).
+Shell Game has a simpler entity-state and action space than Chess.
+However, while many frontier models have been trained on UCI/FEN encoded chess games, the Shell Game is, to the best of our knowledge, not present in the training data of these models, though there may be analogous tasks in the pre-training data.
+We simulate repeated Shell Game swaps to create a set of Shell Game trajectories.
+Swap actions are either: Text-Encoded Swaps (‘‘x swap y’’, where $\{x,y\}\subset\{1,2,3\}$); Image-Encoded Swaps (an image depicting the shells being swapped, with the ball not visible; see Fig.
+1).
 The ground truth entity state after the game finishes is a single number indicating the final shell position of the ball.
-2.3 Models
-We use MET-Bench to evaluate the limitations of frontier models, including vision-language models (VLMs) which accept images and text as input, and newer reasoning models like OpenAI’s o1 that are trained using reinforcement learning and utilize test-time search algorithms to improve their reasoning abilities on domains like mathematics and coding.
-A full list of models and their capabilities is shown in Appendix A, Table 5.
+2.3 Reinforcement Learning for Reasoning
+To improve the entity state reasoning of open-source models, we utilize Group Relative Policy Optimization (GRPO) (shao2024deepseekmath), a policy-gradient RL method for training language models.
+GRPO is suited to tasks with verifiable outcomes, such as entity tracking, where correctness can be determined programmatically.
+By using the task’s ground-truth accuracy as the reward signal, GRPO can directly optimize for entity tracking performance without requiring human preference data or learned reward models.
 3 Methods
@@
 e7e5
-Now what is the final FEN? Only output the FEN.
-Assistant rnbqkbnr/pppp1ppp/8/4p3/4P3/ 8/PPPP1PPP/RNBQKBNR w KQkq - 2 2
+Now what is the final FEN? Output FINAL ANSWER: [FEN].
+Assistant FINAL ANSWER: "rnbqkbnr/pppp1ppp/8/4p3/4P3 /8/PPPP1PPP/RNBQKBNR w KQkq - 0 2"
 Figure 2: An example zero-shot user–assistant exchange in the Chess domain, showing the initial board state as FEN, two UCI moves (e2e4, e7e5) and the final state.
@@
 The FEN is line-broken for readability.
-We utilize the standard chat-based schema exposed by current frontier models that consists of interleaved user-provided and assistant (model) provided messages.
-Figures 2 shows the prompting strategy used for the Chess domain.
-Similarly, Appendix A, 6 shows the prompting strategy used for the Shell Game domain.
-Text actions are represented using simple notation on which the models have been trained, UCI for Chess and a simple domain-specific-language for Shell Game.
+We utilize the standard chat-based input schema exposed by current models that consists of interleaved user-provided and model-generated messages.
+Figure 2 shows the prompting strategy used for the Chess domain.
+Similarly, Appendix A details the prompting strategy used for the Shell Game domain.
 In the case of image-action input, the text actions are replaced with their image-rendered versions as Base64 encoded PNG images and a text description of how to interpret the image-actions is provided.
@@
 Various common image representations were explored including arrows, bounding boxes, and symbolic markers.
-The image depiction of the game actions is explained to the language model every time images are provided using the prompts in Appendix A, Figures 7 & 8.
+The image depiction of the game actions is explained to the language model every time images are provided using the prompts in the Appendix.
+3.1 GRPO Training for Entity Tracking
+To investigate whether reinforcement learning can improve entity tracking in open-weight vision-language models, we apply GRPO to finetune models on MET-Bench tasks.
+Reward Function Design
+For entity tracking, we design reward functions that leverage the automatic verifiability of our benchmark tasks. For Chess, the reward is the fraction of correctly predicted board squares:
+$R_{\text{chess}}(y,y^{*})=\frac{1}{64}\sum_{i=1}^{64}\mathbbm{1}[y_{i}=y^{*}_{i}],$
+where $y$ is the predicted FEN board state and $y^{*}$ is the ground truth. For the Shell Game, the reward is binary based on correct ball position prediction:
+$R_{\text{shell}}(y,y^{*})=\mathbbm{1}[y=y^{*}].$
+We train Gemma3 4B and 27B (gemmateam2025gemma3technicalreport). Full training configurations are in Appendix A.
+(a) Chess accuracy with text actions.
+(b) Chess accuracy with image actions.
+Figure 3: In the text action setting, the reasoning models Claude 3.7 Sonnet Thinking and Gemini 2.5 Pro, maintain the highest accuracy at longer action-sequence lengths.
+All models struggle to maintain accurate board representations in the image action setting, with Claude 3.7 Sonnet Thinking performing the best.
+Bars show 95% confidence intervals.
+(a) Shell Game accuracy with text actions.
+(b) Shell Game accuracy with image actions.
+Figure 4: In the text action setting, the reasoning models Claude 3.7 Sonnet Thinking and Gemini 2.5 Pro achieve the highest performance over long action sequences.
+In the image action setting the accuracy of all models except Gemini 2.5 Pro decreases to random guessing by 2020 actions.
+Bars show 95% confidence intervals.
 4 Experiments
-Chess Text Image
+We perform experiments across a wide range of models and settings to evaluate different aspects of frontier-model entity-tracking performance.
+For all experiments, the models are sampled with a temperature of zero.
+4.1 Tracking in Text Outperforms Images
+We evaluate the difference in accuracy when tracking entity states from text and image actions in the zero-shot, few-shot, chain-of-thought, and reasoning settings and report 95% confidence intervals (CI).
+For the few-shot experiments, $N=5$ examples are selected at random from the training set and prepended to the test example.
+To evaluate the effect of chain-of-thought reasoning, we prompt the model as in the zero-shot case but append “think step by step before producing a final answer.” In both domains, we evaluate on a set of $500$ games selected at random from the test set, each with a sequence length of ten actions, and $100$ games for reasoning models.
+Chess
+Table 1: Entity tracking accuracy (95% CI) in Chess for text-only and image-only actions.
+In the Few‐Shot setting $N=5$ in-context examples are used.
+Methods with explicit reasoning perform best.
+The baseline is predicting the Game Start board configuration.
+Model Chess
+Text Image
 Baseline
-Game Start 74.4 74.4
+Game Start 74.9 ± 0.5 74.9 ± 0.5
 Zero-Shot
-Claude 3.5 Sonnet 96.8 66.2
-MiniMax-VL-01 86.3 65.8
-Gemini-2.0-Flash 93.3 74.7
-GPT-4o mini 68.3 60.6
-GPT-4o 89.6 73.3
+GPT-4o 91.6 ± 0.3 76.0 ± 0.5
+GPT-4o-mini 75.6 ± 0.5 55.3 ± 0.5
+GPT-4.1 85.6 ± 0.4 67.7 ± 0.5
+GPT-4.1-mini 82.5 ± 0.4 77.5 ± 0.5
+GPT-4.1-nano 77.5 ± 0.5 73.3 ± 0.5
+Gemini-2.5-Flash 91.0 ± 0.3 66.9 ± 0.5
+LLaMA-4 Maverick 86.7 ± 0.4 51.1 ± 1.2
+MiniMax-VL-01 85.9 ± 0.4 73.4 ± 0.5
+Claude 3.7 Sonnet 96.1 ± 0.2 70.2 ± 0.5
 Few-Shot (n=5)
-Claude 3.5 Sonnet 99.6 77.9
-MiniMax-VL-01 88.0 75.9
-Gemini-2.0-Flash 96.6 78.7
-GPT-4o mini 75.1 74.9
-GPT-4o 94.9 77.7
+GPT-4o 94.3 ± 0.2 77.6 ± 0.5
+GPT-4o-mini 74.5 ± 0.5 74.2 ± 0.5
+GPT-4.1 86.6 ± 0.4 64.9 ± 0.5
+GPT-4.1-mini 81.2 ± 0.4 76.9 ± 0.5
+GPT-4.1-nano 74.5 ± 0.5 74.6 ± 0.5
+Gemini-2.5-Flash 91.3 ± 0.3 72.0 ± 0.5
+LLaMA-4 Maverick 85.8 ± 0.4 48.1 ± 0.6
+MiniMax-VL-01 88.0 ± 0.8 40.8 ± 1.2
+Claude 3.7 Sonnet 99.2 ± 0.1 77.7 ± 0.5
 Chain-of-Thought
-Claude 3.5 Sonnet 97.9 68.4
-MiniMax-VL-01 62.7 58.0
-Gemini-2.0-Flash 93.3 74.5
-GPT-4o mini 72.0 69.6
-GPT-4o 91.8 74.9
+GPT-4o 94.9 ± 0.2 67.5 ± 0.5
+GPT-4o-mini 68.8 ± 0.5 43.1 ± 0.5
+GPT-4.1 98.1 ± 0.1 75.3 ± 0.5
+GPT-4.1-mini 86.4 ± 0.4 77.6 ± 0.5
+GPT-4.1-nano 49.0 ± 0.6 34.6 ± 0.5
+Gemini-2.5-Flash 77.0 ± 1.0 44.6 ± 1.2
+LLaMA-4 Maverick 82.4 ± 0.4 61.9 ± 0.5
+MiniMax-VL-01 62.3 ± 0.5 32.8 ± 0.5
+Claude 3.7 Sonnet 99.5 ± 0.1 96.2 ± 0.2
 Reasoning
-o1-mini 65.1 -
-o3-mini 99.6 -
-o1 98.2 83.5
-(a) Chess with ten moves in each sequence.
-Shell Game Text Image
+o1 98.2 ± 0.3 83.5 ± 0.9
+o3 99.9 ± 0.1 45.5 ± 1.2
+o4-mini 84.2 ± 0.9 78.0 ± 1.0
+Gemini-2.5-Pro 77.4 ± 1.0 76.8 ± 1.0
+Gemini-2.5-Flash 40.2 ± 1.2 17.4 ± 0.9
+Claude 3.7 Sonnet 99.8 ± 0.1 96.0 ± 0.5
+Claude 4.5 Opus 99.9 $\pm$ 0.1 99.5 $\pm$ 0.2
+GPT-5.2 98.0 $\pm$ 0.4 99.3 $\pm$ 0.2
+In the Chess domain, we report the per-square accuracy of the predicted board, that is the ratio of correctly predicted pieces (or absence of a piece) to the total number of board tiles.
+The ‘Game Start’ baseline is the accuracy of predicting the initial board configuration.
+After only ten actions, much of the board remains unchanged, so this is a strong baseline.
+Table 1 reports the accuracy for text and image actions in the chess domain.
+Performance in the text modality is significantly better than in the image modality.
+In the zero-shot setting, the best-performing text model Claude 3.7 Sonnet, achieves an impressive $96.1\%\pm 0.2\%$ accuracy, while its image-based counterpart drops sharply to $70.2\%\pm 0.5\%$.
+A similar trend holds across models, indicating that current models can perform entity tracking in text but fail to integrate visual updates as effectively.
+Both in-context learning with few-shot, chain-of-thought, and reasoning lead to performance improvements. But the text tracking accuracy exceeds the image accuracy.
+Shell Game
+Table 2: Entity tracking accuracy (95% CI) in Shell Game for text-only and image-only actions. In the Few‐Shot setting $N=5$ in-context examples are used. Methods with explicit reasoning perform best.
+Model Shell
+Text Image
 Baseline
@@
 Zero-Shot
-Claude 3.5 Sonnet 34.4 36.2
-MiniMax-VL-01 34.4 35.2
-Gemini-2.0-Flash 30.0 33.4
-GPT-4o mini 30.6 32.8
-GPT-4o 36.0 32.4
+GPT-4o 33.0 ± 4.1 32.2 ± 4.1
+GPT-4o-mini 33.6 ± 4.1 32.4 ± 4.1
+GPT-4.1 32.2 ± 4.1 36.4 ± 4.2
+GPT-4.1-mini 31.4 ± 4.1 30.8 ± 4.0
+GPT-4.1-nano 31.2 ± 4.0 30.8 ± 4.0
+Gemini-2.5-Flash 35.0 ± 4.2 37.0 ± 4.2
+LLaMA-4 Maverick 30.8 ± 4.0 33.2 ± 4.1
+MiniMax-VL-01 30.8 ± 4.0 31.2 ± 4.0
+Claude 3.7 Sonnet 35.4 ± 4.2 37.8 ± 4.2
 Few-Shot (n=5)
-Claude 3.5 Sonnet 34.0 30.6
-MiniMax-VL-01 36.4 32.0
-Gemini-2.0-Flash 37.0 31.4
-GPT-4o mini 34.4 31.2
-GPT-4o 37.2 31.0
+GPT-4o 33.6 ± 4.1 32.0 ± 4.1
+GPT-4o-mini 34.8 ± 4.2 32.2 ± 4.1
+GPT-4.1 34.6 ± 4.2 33.0 ± 4.1
+GPT-4.1-mini 36.4 ± 4.2 32.8 ± 4.1
+GPT-4.1-nano 36.6 ± 4.2 36.4 ± 4.2
+Gemini-2.5-Flash 31.4 ± 4.1 35.2 ± 4.2
+LLaMA-4 Maverick 35.4 ± 4.2 35.2 ± 4.2
+MiniMax-VL-01 36.4 ± 4.2 32.0 ± 4.1
+Claude 3.7 Sonnet 32.4 ± 4.1 36.0 ± 4.2
 Chain-of-Thought
-Claude 3.5 Sonnet 97.4 94.2
-MiniMax-VL-01 92.6 32.8
-Gemini-2.0-Flash 76.8 33.8
-GPT-4o mini 84.4 35.0
-GPT-4o 99.8 84.2
+GPT-4o 99.0 ± 0.9 35.8 ± 4.2
+GPT-4o-mini 61.0 ± 4.3 30.4 ± 4.0
+GPT-4.1 99.8 ± 0.5 37.6 ± 4.2
+GPT-4.1-mini 100.0 $\pm$ 0.4 72.0 ± 3.9
+GPT-4.1-nano 61.8 ± 4.2 32.2 ± 4.1
+Gemini-2.5-Flash 94.0 ± 4.8 34.0 ± 9.1
+LLaMA-4 Maverick 77.0 ± 3.7 34.6 ± 4.2
+MiniMax-VL-01 77.4 ± 3.7 34.4 ± 4.2
+Claude 3.7 Sonnet 100.0 $\pm$ 0.4 77.4 ± 3.7
 Reasoning
-o1-mini 99.8 -
-o3-mini 100.0 -
-o1 100.0 92.8
-(b) Shell Game with five swap moves in each sequence.
-Table 1: Entity tracking accuracy in Chess and Shell Game for text and image actions.
-In the Few-Shot setting $N=5$ in-context examples are used.
-Methods and models which employ explicit reasoning perform best (chain-of-thought and reasoning models).
-We perform experiments across a wide range of models and settings to evaluate different aspects of frontier-model entity-tracking performance.
-For all experiments, the models are sampled with a temperature of zero.
-4.1 Tracking Entities with Text and Image Actions
-We evaluate difference in accuracy when tracking images from text and image actions in the zero-shot, few-shot, and chain-of-thought settings. These results are presented in Section 5.1.
-Zero-shot
+o1 100.0 $\pm$ 1.9 92.6 ± 2.3
+o3 100.0 $\pm$ 1.9 63.0 ± 9.3
+o4-mini 100.0 $\pm$ 1.9 33.0 ± 4.1
+Gemini-2.5-Pro 100.0 $\pm$ 1.9 100.0 $\pm$ 1.9
+Gemini-2.5-Flash 100.0 $\pm$ 1.9 42.0 ± 4.1
+Claude 3.7 Sonnet 100.0 $\pm$ 1.9 87.0 ± 6.6
+Claude 4.5 Opus 100.0 $\pm$ 1.9 100.0 $\pm$ 1.9
+GPT-5.2 100.0 $\pm$ 1.9 99.0 $\pm$ 2.6
+The final state is a single number $n\in\{1,2,3\}$ that gives the position of the ball, and we measure the accuracy of predicting it. The naive baseline picks a position uniformly at random.
+Table 2 reports the accuracy for text and image actions for Shell Game.
+The results follow a pattern similar to that of Chess, with text-based tracking outperforming image-based tracking.
+However, unlike in Chess, the performance in the zero-shot and few-shot settings are close to random.
+Chain-of-thought and reasoning lead to large performance increases and multiple models attain near perfect accuracy in text.
+These results suggest that when guided to decompose the task step-by-step, models can reason more effectively (but still largely imperfectly) using image inputs, a finding that complements the results of performing cascaded inference as discussed below.
+Table 3: Cascaded entity tracking accuracy for Chess and Shell (Image → Text).
+In cascaded inference, the model is first used to map each image action to the text representation of the action.
+The model is prompted to perform the entity tracking task as in the text-action setting.
+Cascaded
+Method Chess Shell
+Chain-of-Thought
+Gemma3 27B 55.7 63.0
+GPT-4o mini 66.4 47.6
+GPT-4o 93.2 99.4
+4.2 Reasoning Aids Long Sequence Accuracy
+We varied the sequence length of the actions to quantify the effect of compounding errors on the models.
+These sequences range from zero to $100$ actions in both the image and text action modality.
+This serves to quantify the relative drop-off in performance among models in the reasoning and chain-of-thought settings.
 Chess
-In the Chess domain, we evaluate on a set of $100$ games selected at random from the test set, each with a sequence length of ten actions.
-The model must predict the FEN string for the final state.
-If the FEN string contains syntax errors such that it cannot be parsed, the accuracy for that instance is zero.
-We report the per-square accuracy of the predicted board, that is the ratio of correctly predicted pieces (or absence of a piece) to the total number of board tiles.
-The ‘Game Start’ baseline is the accuracy of predicting the initial board configuration.
-After only ten actions, most of the board configuration remains unchanged, so this is a strong baseline.
+Figure 3(a) plots model accuracy against increasing sequence lengths of text actions.
+The models are evaluated in the chain-of-thought setting for sequence lengths of zero to $100$ text actions.
+Reasoning models like Claude 3.7 Sonnet Thinking are able to handle longer sequence lengths with a smaller decrease in accuracy.
+In contrast, the non-reasoning models experience sharp decreases in accuracy after only a few actions.
+Figure 3(b) plots model accuracy against increasing sequence lengths of image actions.
+The models are evaluated in the chain-of-thought setting for sequence lengths of zero to $20$ image actions.
 Shell Game
-We evaluate Shell Game using a set of $500$ games generated at random, each with a sequence length of five actions.
-The final state is a single number $n\in\{1,2,3\}$ that gives the position of the ball, and we measure the accuracy of predicting it.
-The naive baseline picks a position uniformly at random.
-Few-Shot and Chain-of-Thought
-In these settings, the evaluation and procedure remain largely unchanged from the zero-shot Chess.
-For the few-shot experiments, $N=5$ examples are selected at random from the training set and prepended to the test example.
-The few-shot examples have the same number of actions as the test-set examples.
-To evaluate the effect of chain-of-thought reasoning, we prompt the model to ‘think step by step before producing a final answer.’
-4.2 Sequence-Length Variation
-The sequence length of the actions is varied to quantify the effect of compounding errors on the models.
-These sequences range from zero to $100$ actions in both the image and text action modality.
-This serves to quantify the relative drop-off in performance among models in the zero-shot setting.
-These results are presented in Section 5.2.
-4.3 Image-Action Classification
-We perform an experiment to demonstrate that VLMs have the ability to accurately interpret the actions depicted in the image-action representations.
-Figures 7 & 8 in Appendix A show the prompting strategy used to get a VLM to predict (zero-shot) a text action from an image action.
-We perform this evaluation on a test set of $1000$ image depictions of actions for each game.
-The results are presented in Section 5.3.
-4.4 Cascaded Inference
-Using the text actions predicted from the images in the image-action classification task, we devise an ablation to test the effect of cascading (performing multimodal inference in two steps) in the zero-shot setting.
-Given a sequence of image-based observations $\bm{A_{t}^{\text{image}}}$, a vision-language model (VLM) first predicts the corresponding text-based action sequence $\hat{\bm{A}}^{\text{text}}=g(\bm{A}^{\text{image}})$, where $g$ maps images to text using the procedure in Section 4.3.
+In the text modality in Figure 4(a), the reasoning models Gemini 2.5 Pro and Claude 3.7 Sonnet Thinking attain the highest accuracies.
+Gemini 2.5 Pro performs nearly perfectly at a sequence length of $100$ actions, where the non-reasoning models’ performance is significantly degraded.
+In the image modality results in Figure 4(b), Gemini 2.5 Pro performs better than the other models, which see a rapid decrease in accuracy with sequence lengths longer than ten actions.
+By $20$ actions, the performance of all models except Gemini 2.5 Pro has degraded to random guessing.
+The superior performance of reasoning models suggests that structured inference mechanisms, such as test-time search or chain-of-thought, are crucial for maintaining coherent entity states over long sequences.
+This aligns with prior findings that models trained on structured reasoning tasks (e.g., mathematics, coding) develop stronger implicit state-tracking capabilities (kim2024codepretrainingimprovesentity), whereas standard vision-language models struggle to track entities beyond short visual action sequences.
+Table 4: Percent image‐action classification accuracy (with 95% CI) for various models.
+We report the accuracy of predicting the shell swap/UCI action for both Chess and Shell Game on 1,0001{,}000 image actions.
+Model Perception Accuracy (%)
+Chess
+GPT-4o-mini 46.4 ± 1.0
+GPT-4o 94.5 ± 1.4
+Claude 4.5 Opus 100.0 ± 0.2
+Gemma3 27B 67.5 ± 2.9
+Shell Game
+GPT-4o-mini 100.0 ± 0.2
+GPT-4o 100.0 ± 0.2
+Claude 4.5 Opus 100.0 ± 0.2
+Gemma3 27B 100.0 ± 0.2
+4.3 Models Understand Image Actions
+Table 4 shows the performance of models on predicting the text action represented by each image.
+While some models fail to accurately localize the Chess moves, the evaluated models understand the Shell Game actions with high accuracy.
+This indicates that perception of the image-actions is not the only limiting factor for effective entity tracking with image inputs.
+4.4 Cascading Matches Text-Only Tracking
+To differentiate the effects of perception and visual reasoning, we evaluate cascaded inference.
+The performance in this cascaded setting is similar to the text-action performance, showing that the model has the task-knowledge needed to perform entity tracking in both domains, but cannot reason effectively directly from the image modality.
+Using the text actions predicted from the images in the image-action classification task, we devise a cascaded approach to test the effect of performing multimodal inference in two separate steps.
+Given a sequence of image-based observations $\bm{a_{t}^{\text{image}}}$, a vision-language model (VLM) first predicts the corresponding text-based action sequence $\hat{\bm{a}}^{\text{text}}=g(\bm{a}^{\text{image}})$, where $g$ maps images to text using the procedure described in the previous section.
 The text actions are then used for zero-shot inference to estimate the final state as $\bm{S_{T}}=f(\bm{S_{0}},\hat{\bm{a}}^{\text{text}})$.
 This removes the need for the model to perform entity tracking directly in the image modality, isolating the effect of perceptual failures.
-These results are presented in Section 5.4.
-4.5 Fine-Tuning Experiments
-We fine-tune frontier OpenAI models using their fine-tuning API on a small training set.
-We evaluate these in the Chess and Shell Game domains using both image and text modalities.
-For Chess, we select a random set of $100$ training examples of sequence length ten.
-For Shell Game we use $20$ training examples of sequence length $3$.
-Results are available in Section A.2.
-Hyperparameters were determined through grid search and are available in Appendix A.2, Table 6.
-4.6 Mixed-Modality Experiments
-We perform an ablation where the image-action and text-action modalities are mixed.
-At each action step, one action modality is randomly selected with a probability varying from $100\%$ text actions to $100\%$ image actions.
-This serves to quantify the effect of forcing the model to reason over multiple input modalities simultaneously.
-These results are presented in Section A.1.
-5 Results
-5.1 Tracking in Text Outperforms Images
+Table 3 shows results for this approach in the zero-shot setting.
+The image actions are first explicitly translated into text actions, and then run through the text-based entity tracking pipeline.
+4.5 GRPO Improves Entity Tracking
+We train open-source VLMs for improving entity tracking on a 5-action version of MET-Bench.
+GRPO is highly effective at teaching entity tracking to models, achieving performance on par with several advanced closed models.
+However, the lack of cross-modal transfer suggests models process entity tracking in text and image through different pathways.
+Robust multimodal entity tracking may require architectural innovations or training strategies beyond what RL provides.
+Table 5: Entity tracking accuracy (%) after RL training for Gemma3 4B and 27B models. GPT-4o-mini model for comparison. 95% confidence intervals.
+Model Text Image
+Shell Game
+Gemma3-4B (Base) 44.4 $\pm$ 4.3 32.2 $\pm$ 4.1
+Gemma3-4B (GRPO Text) 71.0 $\pm$ 4.0 32.2 $\pm$ 4.1
+Gemma3-27B (Base) 80.4 $\pm$ 3.5 44.4 $\pm$ 4.3
+Gemma3-27B (GRPO Image) 79.8 $\pm$ 3.5 62.0 $\pm$ 4.2
+GPT-4o-mini 86.4 $\pm$ 3.0 36.8 $\pm$ 4.2
 Chess
-Table 1(a) reports the accuracy for text and image actions in the chess domain.
-Performance in the text modality is significantly better than in the image modality.
-In the zero-shot setting, the best-performing text model, Claude 3.5 Sonnet, achieves an impressive $96.8\%$ accuracy, while its image-based counterpart drops sharply to $66.2\%$.
-A similar trend holds across models, indicating that current models can perform entity tracking in text but fail to integrate visual updates as effectively.
-Both in-context learning with few-shot prompting and chain-of-thought reasoning lead to performance improvements.
-In the image modality, the accuracy is not significantly greater than the naive baseline.
-Only the o1 reasoning model achieves an accuracy greater than $80\%$ in the image modality.
-Shell Game
-(a) Chess zero-shot accuracy with text actions.
-(b) Chess zero-shot accuracy with image actions.
-Figure 3: In the text action setting, the reasoning models, o1 and o3-mini, maintain the highest accuracy at longer sequence lengths.
-o1-mini and the other models begin to output invalid and inaccurate board representations.
-All models struggle to maintain accurate board representations in the image action setting, with o1 performing the best.
-Table 1(b) reports the accuracy for text and image actions for Shell Game.
-The results follow a pattern similar to that of Chess, with text-based tracking outperforming image-based tracking.
-However, unlike in Chess, the performance in the zero-shot setting is close to random.
-But the best model, o1, attains accuracies of $100.0\%$ and $92.8\%$ for the text and image modalities, respectively.
-Few-shot prompting provides only marginal improvements, but chain-of-thought gives large performance increases.
-GPT-4o’s accuracy jumps from $36.0\%$ to $99.8\%$ in text, and from $32.4\%$ to $84.2\%$ in image tracking.
-These results suggest that when guided to decompose the task step-by-step, models can reason more effectively using image inputs, a finding that complements the results of performing cascaded inference in Section 5.4.
-5.2 Reasoning Aids Long Sequence Accuracy
-Chess
-Figure 3(a) plots model accuracy against increasing sequence lengths of text actions.
-The models are evaluated in the zero-shot setting for sequence lengths of zero to $100$ text actions.
-Reasoning models like o1 and o3-mini are able to handle longer sequence lengths with a smaller decrease in accuracy.
-However, o1-mini performs worse than the other models as it produces more invalid FEN board representations at longer sequence lengths.
-In contrast, the non-reasoning models experience sharp decreases in accuracy after only a few actions.
-Figure 3(b) plots model accuracy against increasing sequence lengths of image actions.
-The models are evaluated in the zero-shot setting for sequence lengths of zero to $20$ image actions.
-While the reasoning models attain higher accuracies, the performance differences are smaller than in the text-action setting.
-Shell Game
-In the text modality in Figure 4(a), the reasoning models o1, o1-mini, and o3-mini attain the highest accuracies.
-o1 performs perfectly at a sequence length of 50 actions, where the non-reasoning models’ performance is significantly degraded.
-In the image modality results in Figure 4(b), o1 performs better than the other models, but sees a rapid decrease in accuracy with sequence lengths longer than five actions.
-By $20$ actions, the performance of all models has degraded to random.
-The superior performance of reasoning models like o1 suggests that structured inference mechanisms, such as test-time search or chain-of-thought, are crucial for maintaining coherent entity states over long sequences.
-This aligns with prior findings that models trained on structured reasoning tasks (e.g., mathematics, coding) develop stronger implicit state-tracking capabilities (Kim et al., 2024), whereas standard vision-language models struggle with entity persistence beyond short contexts.
-5.3 Models Understand Image Actions
-(a) Shell Game zero-shot accuracy with text actions.
-(b) Shell Game zero-shot accuracy with image actions.
-Figure 4: In the text action setting, the reasoning models o1, o1-mini, and o3-mini achieve the highest performance over long action sequences, but performance degrades for all models as sequence length increases.
-In the image action setting, o1 performs the best, but achieves an accuracy no better than guessing the starting state by $20$ actions.
-Model Start End Overall
-Chess
-GPT-4o-mini 62.1 55.2 41.0
-GPT-4o 97.3 97.0 94.5
-Shell Game
-GPT-4o-mini 100.0 100.0 100.0
-GPT-4o 100.0 100.0 100.0
-Table 2: Percent image-action classification accuracy for various models.
-We report the accuracy of predicting the action start, end, and overall/UCI action for both Chess and Shell Game on $1{,}000$ image actions.
-Cascaded
-Method Chess Shell
-Baseline
-Game Start 74.4 33.3
-Zero-Shot
-GPT-4o mini 63.0 28.0
-GPT-4o 89.8 34.0
-Table 3: Cascaded entity tracking accuracy for Chess and Shell (Image → Text).
-In cascaded inference, the model is first used to map each image action to the text representation of the action.
-Then model is prompted to perform the entity tracking task as in the text-action setting.
-Table 2 shows the performance of GPT-4o and GPT-4o-mini on classifying the text action represented by each image action.
-For Chess, the action ‘start’ is the square of the piece being moved, and ‘end’ is the destination square.
-For Shell Game, the ‘start’ is the first shell to be swapped and ‘end’ is the second.
-‘Overall’ is the accuracy of classifying the entire action (start and end) correctly.
-While GPT-4o-mini struggles to recognize actions in chess, GPT-4o achieves an accuracy of $94.5\%$.
-Both models attain perfect accuracy on the simpler Shell Game domain.
-This indicates that perception of the image-actions is not the fundamental limiting factor for effective entity tracking with image inputs.
-A potential concern is whether the observed failures stem from poor image representation rather than reasoning deficiencies.
-However, our cascaded inference experiments (Section 5.4) demonstrate that when models first translate image actions into text, they achieve near-text-level accuracy.
-This suggests that models can correctly parse image-based actions but struggle with integrating them into coherent state updates, a limitation in reasoning rather than perception.
-5.4 Cascading Matches Text-Only Tracking
-Table 3 shows the accuracy of cascaded inference in the zero-shot setting for GPT-4o and GPT-4o-mini.
-In this setting, the image actions are first translated into text actions, and then run through the text-based entity tracking pipeline.
-The performance in the cascaded setting is similar to the text-action performance, showing that the model has the task-knowledge needed to perform entity tracking in both domains, but cannot reason effectively in the image modality.
+Gemma3-4B (Base) 66.1 $\pm$ 1.2 14.6 $\pm$ 0.9
+Gemma3-4B (GRPO Text) 84.4 $\pm$ 0.9 26.3 $\pm$ 1.1
+Gemma3-27B (Base) 75.8 $\pm$ 1.0 63.1 $\pm$ 1.2
+Gemma3-27B (GRPO Image) 73.8 $\pm$ 1.1 74.5 $\pm$ 1.1
+GPT-4o-mini 77.0 $\pm$ 0.5 60.1 $\pm$ 0.5
+Table 5 shows the results of GRPO training on the Shell Game domain.
+When trained on text-only inputs with chain-of-thought prompting, Gemma3-4B improves from 44.4% (near random) to 71.0% accuracy.
+Figure 5 illustrates the learning dynamics.
+Limited Cross-Modal Transfer
+Despite the improvement in single modality entity tracking, RL reasoning training does not transfer robustly to the held-out modality.
+In Shell Game, the held-out modality does not improve at all.
+This finding suggests that the entity tracking capabilities learned through text-based RL training are modality-specific and do not generalize to other modalities.
+The apparent transfer in Chess image modality can be attributed to the model learning to output more valid FENs.
+Shell, Image Actions (Gemma3 27B) Shell, Text Actions (Gemma3 4B)
+Refer to caption Refer to caption
+Chess, Image Actions (Gemma3 27B) Chess, Text Actions (Gemma3 4B)
+Refer to caption Refer to caption
+Figure 5: GRPO training curves across domains and modalities.
+Each series shows test accuracy for the train modality and transfer modality for image ($\bullet$) and text ($\blacksquare$).
+GRPO improves in-modality performance but this does not substantially transfer to the other modality.
+95% confidence intervals.
 5 Discussion
-Our evaluation of frontier model performance on MET-Bench provides several insights into the current state and remaining challenges of multimodal entity tracking.
-We demonstrate a significant performance gap between text-based and image-based entity tracking across all evaluated models, with even state-of-the-art vision-language-reasoning models struggling to maintain accurate entity states when processing visual inputs.
-This disparity persists across both the Chess and Shell Game domains, suggesting a fundamental limitation in current architectures’ ability to reason about entity states through visual observations.
-This finding is particularly noteworthy given that our image-action classification results (Table 2) demonstrate that models can accurately perceive and classify individual visual actions.
+Many VLM applications rely on robust multimodal entity tracking.
+Our evaluation of frontier models on MET-Bench provides several insights into the current remaining challenges of multimodal entity tracking.
+We demonstrate a significant performance gap between text-based and image-based entity tracking across many state-of-the-art vision-language-reasoning models.
+This disparity generally persists across both the Chess and Shell Game domains, suggesting a fundamental limitation in current architectures’ ability to reason about entity states through visual observations.
+This finding is particularly noteworthy given that our image-action classification results (Table 4) demonstrate that models can accurately perceive and classify individual visual actions in Shell Game.
 The gap between perception and reasoning suggests that the challenge lies not in processing visual inputs, but in maintaining and updating coherent entity information across sequential visual observations.
 Our cascaded inference experiments provide further evidence for this interpretation.
-When models first translate visual inputs to text before performing entity tracking, they achieve performance comparable to pure text-based tracking.
+In Shell Game, when models first translate visual inputs to text before performing entity tracking, they achieve performance comparable to pure text-based tracking.
 This indicates that the models possess the relevant task knowledge and reasoning capabilities, but struggle to apply them directly in the visual domain.
-Further, the effectiveness of chain-of-thought prompting, particularly in the Shell Game domain where it improved GPT-4o’s accuracy from $36.0\%$ to $99.8\%$ for text and $32.4\%$ to $84.2\%$ for images, highlights the importance of explicit reasoning for entity tracking.
-This improvement indicates that current models can perform complex entity tracking when guided to decompose the task into smaller steps, even in novel domains not present in their training data.
-However, the fact that such prompting was necessary suggests that models do not implement robust tracking, particularly in multimodal settings.
-Lastly, the performance of specialized reasoning models like o1 and o3-mini on longer sequences demonstrates the potential of architectures explicitly trained for sequential reasoning to maintain coherent entity states despite the challenges of accumulating errors over extended sequences.
+Further, the effectiveness of chain-of-thought prompting, particularly in the Shell Game domain where it improved Claude 3.7 Sonnet’s accuracy from near random to $100.0\%$ for text and $77.4\%$ for images, highlights the importance of explicit reasoning for entity tracking.
+Current models can perform complex entity tracking when guided to decompose the task into smaller steps, even in novel domains not explicitly present in their training data.
+The RL experiments demonstrate that reasoning training can substantially improve entity tracking within a single modality, with Gemma3-4B improving from near-random performance to 71.0% accuracy on text-based Shell Game tracking.
+However, the lack of robust cross-modal transfer suggests that current VLM architectures learn modality-specific representations rather than unified entity tracking capabilities.
+Achieving robust multimodal entity tracking may require new methods, such as explicit cross-modal alignment objectives or other inductive biases that encourage models to develop modality-agnostic tracking mechanisms.
 6 Related Work
-Entity tracking has been extensively studied in textual domains, with a focus on probing and improving language models’ abilities to maintain representations of entity states.
-For instance, Toshniwal et al.
-(2022) evaluates chess as an entity tracking domain, employing fine-tuned models (Radford et al., 2019) to assess performance.
-Similarly, Kim and Schuster (2023) examine the impact of model size and fine-tuning on entity tracking in textual settings similar to our Shell Game domain.
+Entity tracking has been extensively studied in textual domains, with a focus on probing and improving LLMs’ abilities to maintain representations of entity states.
+toshniwal2022chess evaluates chess as an entity tracking domain, employing finetuned models (radford2019language) to assess performance.
+Kim and Schuster (2023) examine the impact of model size and fine-tuning on entity tracking in textual settings similar to our Shell Game domain.
 Tandon et al.
@@
 In contrast, our work requires predicting entity state changes from actions specified in images and involves larger state spaces.
-Several studies explore the implicit representations of entity states in language models.
+Several studies explore the implicit representations of entity states in LLMs.
+li2025language employ interpretability methods to understand how LLMs implement entity tracking.
 Li et al.
 (2021) and Long et al.
-(2016) use semantic probing to reveal that Transformer-based models (Vaswani et al., 2017) capture entity state representations implicitly during textual reasoning.
-Building on this, Prakash et al.
-(2024) demonstrate that fine-tuning language models for entity tracking tasks enhances pre-existing internal mechanisms rather than learning entirely new representations.
-Li et al.
-(2023) find that Transformers trained on Othello games form internal representations of the game state.
+(2016) use semantic probing to reveal that Transformer-based models (vaswani2017attention) capture entity state representations implicitly during textual reasoning.
+Building on this, prakash2023fine demonstrate that fine-tuning language models for entity tracking tasks enhances pre-existing internal mechanisms rather than learning entirely new representations.
+li2023emergent find that Transformers trained on Othello games form internal representations of the game state.
 Efforts to improve textual entity tracking beyond domain-specific fine-tuning include Fagnou et al.
@@
 They propose a novel attention mechanism to enhance entity tracking in Transformers.
-Gupta and Durrett (2019) fine-tunes small Transformer-based models for tracking entity state in instructional texts.
-Kim et al.
-(2024) investigates how code pretraining improves language models’ abilities to track entities in text, while Yoneda et al.
-(2024) introduce Stalter, a prompting method designed to maintain accurate state representations in text-based robotics planning.
-These works focus on entity tracking as a unimodal, text-based reasoning task.
-While unimodal approaches have achieved substantial progress, there remains a gap in evaluating models’ ability to integrate multimodal inputs for entity tracking.
-Our work extends these evaluations to the multimodal setting and quantifies the performance improvement of reasoning models for entity tracking.
+Gupta and Durrett (2019) trains small Transformer-based models for tracking entity state in instructional texts.
+kim2024codepretrainingimprovesentity investigates how code pretraining improves language models’ abilities to track entities in text, while yoneda2024statler introduce Statler, a prompting method designed to maintain accurate state representations in text-based robotics planning.
+Recent work demonstrates the effectiveness of reinforcement learning for improving reasoning capabilities in language models.
+DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability) shows that RL training with verifiable rewards can significantly enhance mathematical and logical reasoning with Group Relative Policy Optimization (GRPO) (shao2024deepseekmath).
+We apply GRPO to multimodal entity tracking, demonstrating both its promise for teaching visual reasoning about state updates and its limitations in enabling cross-modal transfer.
 7 Conclusion
-Our findings suggest that the primary bottleneck in multimodal entity tracking is not visual recognition but sequential reasoning over visual updates.
-Unlike text-based representations, which align with the models’ training paradigms, visual updates require implicit state reconstruction—a task that current architectures do not perform reliably.
-Future work should explore the effect of additional visual-reasoning post-training, explicit memory structures, or hybrid symbolic representations to mitigate this gap.
-Additional research directions include investigating the role of entity tracking in world-modeling, narrative understanding, and expanding MET-Bench to include more complex domains beyond games.
-We believe addressing these challenges will be crucial for developing AI systems capable of robust reasoning for real-world tasks.
-9 Limitations
-Our benchmark is restricted to two synthetic domains—Chess and the Shell Game.
-While these domains are well-structured and offer a clear framework for assessing entity state tracking, they may not fully capture the complexity of real-world multimodal reasoning tasks that require entity tracking.
-Our study highlights a substantial performance gap between text-based and image-based entity tracking, yet the exact causes of this disparity remain unclear.
-While our cascaded inference experiments suggest that models struggle to reason directly over visual updates rather than simply perceiving them, further investigation is needed to pinpoint the underlying source of the problem and whether it’s possible to explicitly ameliorate it.
-Additionally, our fine-tuning experiments demonstrate that performance can be improved with domain-specific adaptation, but we do not investigate the trade-offs between fine-tuning and generalization across unseen multimodal datasets.
-Understanding whether these improvements transfer to novel multimodal reasoning tasks is an important avenue for future work.
-Ethical Considerations
-Vision-language models may inherit biases from their training data or other aspects of their development, leading to disparities in performance across different languages, cultures, or other categories that intersect with textual and visual representations.
-The evaluations in this work can be used to improve the capabilities of frontier models, which if deployed in contexts requiring multimodal entity tracking, could further expose end-users to the biases present in this work including our choice of game domains and language.
-Acknowledgments
-This paper and associated code were created with the assistance of ChatGPT for proofreading and debugging.
+Our findings suggest that the primary bottleneck in multimodal entity tracking may not be visual recognition but sequential reasoning over visual updates.
+Unlike text-based representations, which align with the models’ training paradigms, visual updates require implicit state reconstruction, a task that most current architectures do not perform reliably.
+Our RL experiments reveal a promising path forward that can effectively train open-source models, achieving performance comparable to closed models.
+However, the lack of cross-modal transfer underscores that visual entity tracking requires fundamentally different representations or training objectives.
+Future work should explore deficits in entity tracking using mechanistic interpretability methods.
+Additional research directions include novel multimodal representation learning methods and explicit memory structures to mitigate this gap.
+We believe addressing these challenges will be crucial for developing AI systems capable of robust multimodal reasoning for real-world tasks.
+Impact Statement
+As with any work that improves AI reasoning capabilities, advancements in multimodal entity tracking could eventually contribute to more capable autonomous systems.
+While our benchmark focuses on structured game domains (Chess and Shell Game), the underlying capabilities are relevant to robotics and other systems employing automated decision-making.
+We encourage researchers building on this work to consider appropriate safeguards in consequential real-world settings.
 Appendix A Appendix
@@
 However for Shell Game, the opposite is true and mixes of text and image actions are challenging for the model to reason over.
-A.2 Fine-Tuning Improves Multimodal Entity Tracking
-Fine-tuned Text Image
-Chess
-GPT-4o mini 89.2 -
-GPT-4o 97.0 86.4
-Shell Game (S=3)
-GPT-4o mini 32.0 -
-GPT-4o 74.0 32.0
-Table 4: Fine-tuned model entity tracking accuracy for the Chess and Shell Game domains (Text actions and Image actions).GPT-4o-mini does not support image finetuning.
-A training set of $100$ action sequences of length ten were used for Chess and $20$ action sequences of length three and five for Shell Game.
-Fine-tuning using the OpenAI fine-tuning API substantially improves model performance across both text and image modalities, as shown in Table 4.
-In Chess, fine-tuned models outperform even the strongest zero-shot reasoning models, achieving 97.0% accuracy in the text domain and a significant boost to 86.4% in the image domain.
-This suggests that even with a relatively small dataset, fine-tuning allows the model to learn entity tracking representations that generalize better in both modalities.
-Notably, fine-tuning leads to a larger improvement in the image modality than in the text modality.
-This reinforces the idea that pretrained models already encode strong textual reasoning capabilities, whereas multimodal reasoning requires additional adaptation.
-In contrast, improvements on a simplified version of Shell Game with only three moves are minimal in case of image-encoded actions.
-The Shell Game task is not present in the training data and it’s harder for the model to generalize, even when exposed to a large fraction of the possible games of the given length.
-This may indicate that the Shell Game domain is simply too challenging for the model to learn in both the image and text settings from a limited number of examples.
-A more complex training curriculum involving fine-tuning over multiple game lengths may be required.
-Model Name Image Reasoning
-Claude 3.5 Sonnet
-Anthropic (2024)
-Gemini-2.0-Flash
-Hassabis and Kavukcuoglu (2024)
-GPT-4o mini
-OpenAI (2024a)
-GPT-4o
-OpenAI (2024b)
-Minimax-VL-01
-MiniMax et al. (2025)
-o1-mini
-OpenAI (2024c)
-o1
-OpenAI (2024d)
-o3-mini
-OpenAI (2025)
-Table 5: Comparison of capabilities of language models evaluated using the MET benchmark. All evaluated models support text input and output. The total API cost of experiments run is \$$2340.00$.
-Configuration Epochs LRM Batch
-Chess
-   Text:
-    GPT-4o-mini 3 1.8 1
-    GPT-4o 3 2 1
-   Image:
-    GPT-4o 3 2 1
-Shell Game
-   Text:
-    GPT-4o-mini 3 1.8 1
-    GPT-4o 3 2 1
-   Image:
-    GPT-4o 5 2 1
-Table 6: Hyperparameters used for fine-tuning across domains and modalities. The training epochs, learning rate multiplier (LRM), and batch size are reported.
 Role Messages
@@
 Figure 9: An example user–assistant exchange in the Shell Game domain, where the assistant identifies the shell swap represented in the image.
-A.3 Models
-The models evaluated using MET-Bench are listed in Table 5.
-Minimax-VL-01
-This model is released under the license: https://github.com/MiniMax-AI/MiniMax-01/blob/main/LICENSE.
-The model is 465 billion parameters and is trained on a “diverse [dataset] incorporating diverse sources including academic literature, books, web content, and programming code” and post-training dataset encompassing many multimodal and NLP tasks of 512 billion tokens (MiniMax et al., 2025).
-A.3.1 Proprietary Models
-These models have limited information about their training and development.
-Like Minimax-VL-01, these models are likely trained on diverse, web-scale corpora spanning many domains and tasks.
-We provide links to the current terms of their use.
-Claude 3.5 Sonnet
-Gemini-2.0-Flash
-GPT-4o mini, GPT-4o, o1, o1-mini, o3-mini
-A.4 Datasets
-The Chess dataset is adapted from Toshniwal et al.
-(2022) which is adapted from the MillionBase dataset, available for download at https://rebel13.nl/rebel13/rebel%2013.html.
-To the best of our knowledge, no license or terms of use are currently listed for either the original MillionBase dataset or dataset of Toshniwal et al.
-(2022).
-Our usage of this dataset is consistent with the description of its use by Toshniwal et al.
-(2022).
-MET-Bench is intended for evaluating and improving the ability of VLMs to perform entity tracking. It is released under the MIT License.
+A.2 GRPO Training Hyperparameters
+Table 6 summarizes the key hyperparameters used for GRPO training across domains and modalities.
+All experiments are implemented using the TRL library (vonwerra2020trl) with the AdamW optimizer.
+For the 27B model, we employ LoRA (hu2021loralowrankadaptationlarge) with rank 16 to reduce memory requirements.
+All experiments use chain-of-thought prompting and are trained on 100 examples with 5 actions per sequence.
+Table 6: GRPO training hyperparameters for the experiments shown in Figure 5.
+Gemma3-4B Gemma3-27B
+Param Chess Shell Chess Shell
+(Text) (Text) (Image) (Image)
+Optimizer AdamW AdamW AdamW AdamW
+Learning Rate 1e-6 1e-6 1e-5 1e-5
+Batch Size 4 4 1 1
+Grad. Accum. 16 16 16 16
+Group Size 8 8 8 8
+LoRA r=16 r=16